OpenStudy (anonymous):

How do you find the union of two sets? Examples: {2,10,5,9} ^ {5, 11, 2} and {1,10} ^ {1,5,9}

OpenStudy (anonymous):

I think you have indicated the intersection. Are you sure you want the union?

OpenStudy (anonymous):

Well, it asks for the union for these two. What exactly is intersection and how do you write out the answer for both?

OpenStudy (anonymous):

first intersection(take common elements from each pair you indicated and put in a set with out repetition ....2,5,1(Ans)

OpenStudy (anonymous):

Okay, then what?

OpenStudy (anonymous):

next union( put them the common element in each pair without repetition)

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

first get the common elements in each pair then the union is the set which contains those elements(commom elements) with out repetition..is it clear!

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

\[(\left\{ 2,10,5,9 \right\} \bigcap \left\{ 5, 11, 2 \right\})\bigcup( \left\{ 1,10 \right\} \bigcap \left\{ 1,5,9 \right\})\]\[=\left\{ 2,5 \right\}\bigcup \left\{ 1 \right\}=\left\{ 1,2,5 \right\}\]

OpenStudy (anonymous):

Mine is a little sparse on explanation, but follows the instruction above.

OpenStudy (anonymous):

Was that {1,2,5} the union?